# Dynamic Conditional Optimal Transport through Simulation-Free Flows

### v1

@@ Abstract
We study the geometry of conditional optimal transport (COT) and prove a dynamical formulation which generalizes the Benamou-Brenier Theorem. With these tools, we propose a simulation-free flow-based method for conditional generative modeling. Our method couples an arbitrary source distribution to a specified target distribution through a triangular COT plan. We build on the framework of flow matching to train a conditional generative model by approximating the geodesic path of measures induced by this COT plan. Our theory and methods are applicable in the infinite-dimensional setting, making them well suited for inverse problems. Empirically, we demonstrate our proposed method on two image-to-image translation tasks and an infinite-dimensional Bayesian inverse problem.
@@ 1 Introduction
Many fundamental tasks in machine learning and statistics may be posed as estimating or sampling from a distribution, where one only has access to the distribution via a finite set of observations. Such tasks range from generative modeling and density estimation to Bayesian inference. A broad range of methods for sampling from an intractable distribution $p$ may be seen through the lens of measure transport. Intuitively, one fixes a reference distribution $q$ which is easy to sample from, and samples $z\sim q$ are transformed by a learned mapping $T$ such that $T(z)$ is approximately distributed as $p$. The distribution obtained in this way from $q$, denoted $T_{\#}q$, is the pushforward of $q$ along $T$.
Generative models (Ho et al., 2020; Song et al., 2020; Kingma and Welling, 2013; Goodfellow et al., 2014; Papamakarios et al., 2021) typically find such a mapping $T$ by minimizing a divergence between the target and generated distribution. In the standard maximum-likelihood based setting, the mapping $T$ is learned by minimizing (an upper bound on) the KL divergence between $p$ and the pushforward $T_{\#}q$. However, a natural alternative approach, based on optimal transport (Villani et al., 2009; Santambrogio, 2015), is to search for a map $T$ which satisfies the constraint $T_{\#}q=p$ and has minimal cost (in some sense).
In many scenarios, we are interested in conditional sampling. That is, we have access to some observation $y\in Y$, and from this information we would like to infer information regarding an unknown quantity $u\in U$. For instance, in Bayesian inference, $y$ denotes the data and $u$ the parameters of a model. A joint distribution $p(y,u)=p(y\mid u)p(u)$ is specified via a likelihood $p(y\mid u)$ and a prior $p(u)$, and one would like to draw samples from the posterior $p(u\mid y)$. However, in applications such as Bayesian inverse problems (Dashti and Stuart, 2013), one is interested in approximating $p(u\mid y)$ for many different observations $y$. Moreover, in applications such as simulation-based inference (Cranmer et al., 2020), the likelihood function may be intractable while we are still able to sample from $p(y,u)$. This motivates an amortized and likelihood-free (Amos et al., 2023) approach, where a single model may be trained and applied across different observations.
In the framework of measure transport, we first specify a source distribution $q(u)$. For any given, fixed $y\in Y$, we seek to find a mapping $T(y)$, depending on $y$, such that $T(y)_{\#}q(u)\approx p(u\mid y)$. Finding such a mapping for each individual $y$ via optimal transport, though, is typically intractable, as we frequently do not have samples from the posterior $p(u\mid y)$. As such, we must leverage information from other observations. One way to achieve this is through triangular mappings (Baptista et al., 2020), where a joint source distribution $q(y,u)$ is transformed by a mapping of the form $T:(y,u)\mapsto(T_{Y}(y),T_{U}(y,u))$. It can be shown (see Section 3.2) that if $T_{\#}q=p$, then $T_{U}(y,{-})$ almost surely couples $q(u\mid y)$ and $p(u\mid y)$. Thus, if we are able to find such a triangular $T$ using samples from the joint distribution $p(y,u)$, then we may obtain conditional samplers.
In this work, we develop a dynamical framework for conditional optimal transport (COT). In particular, we study absolutely curves of measures in a conditional Wasserstein space over $Y\times U$. Our main results (Theorem 14 and Theorem 15) show that such curves are, informally stated, generated by the flow induced by a triangular vector field. Intuitively, a triangular vector field on $Y\times U$ has no velocity in the $Y$ component and thus induces a flow for each conditional measure. As a consequence, in Theorem 16 we obtain a conditional version of the well-known Benamou-Brenier theorem (Benamou and Brenier, 2000). Importantly, our framework is applicable in infinite-dimensional spaces, making it well suited for function space inference problems.
We apply our results on dynamical COT to design conditional generative models, based on the flow matching framework (Lipman et al., 2022; Albergo et al., 2023a) . In particular, we generalize the methods of Pooladian et al. (2023a) and Tong et al. (2023), which utilize optimal transport to pair source and target samples, to the conditional setting. Our method works by solving a relaxed optimal transport problem (Hosseini et al., 2023) during training. In this way, we amortize solutions to the dynamic COT problem across arbitrary $y$. Our method may use arbitrary source distributions, and is applicable across a wide range of conditional generative modeling tasks.
Specifically, the main contributions of our work are as follows.
@@
    In Section 4, we study the conditional Wasserstein space $\mathbb{P}_{p}^{\mu}(Y\times U)$ equipped with the conditional Wasserstein distance $W_{p}^{\mu}(Y\times U)$. We show that this space is a metric space, and moreover, admits constant speed geodesics between any two points.
@@
    In Section 5, we characterize the absolutely continuous curves of measures in $\mathbb{P}_{p}^{\mu}(Y\times U)$ via the continuity equation and triangular vector fields. As a corollary, we obtain a conditional version of the Benamou-Brenier theorem (Benamou and Brenier, 2000).
@@
    In Section 6 we propose a method for learning dynamic COT maps based on flow matching. In Section 7, we demonstrate our method on two image-to-image translation tasks, as well as an infinite-dimensional Bayesian inverse problem.
@@ 2 Related Work
Optimal Transport.
Machine learning based approaches to unconditional optimal transport plans typically focus on the static problem (Makkuva et al., 2020; Korotin et al., 2022; Taghvaei and Jalali, 2019). Approaches to learning optimal transport maps in a dynamic fashion often require restricted architectures (Bunne et al., 2022b) or use likelihood-based approaches (Tong et al., 2020; Finlay et al., 2020; Onken et al., 2021; Huguet et al., 2022), which require simulating from the model during training. While most work focuses on the squared norm cost, the recent approaches of Pooladian et al. (2023b) and Neklyudov et al. (2023) study dynamical optimal transport with costs specified by a Lagrangian. These frameworks, however, typically have no clear interpretation as a static problem.
Conditional optimal transport (COT), on the other hand, is relatively unexplored. Bunne et al. (2022a) and Wang et al. (2023) learn supervised COT maps in a static setting via convex networks (Makkuva et al., 2020). Hosseini et al. (2023) recently develop a framework for conditional optimal transport that is applicable in infinite-dimensional spaces. This builds on the work of Carlier et al. (2016) which applies COT for vector quantile regression. A common approach to COT is to employ triangular plans (Baptista et al., 2020). Our work develops a dynamic formulation of the static COT problem (Hosseini et al., 2023; Carlier et al., 2016) using a notion of triangular vector fields. Additionally, Wang et al. (2023) propose a likelihood-based approach using continuous normalizing flows, but do not develop the theory of dynamical COT as we do here.
@@
Flow matching (Lipman et al., 2022) and the equivalent stochastic interpolant framework (Albergo and Vanden-Eijnden, 2022; Albergo et al., 2023a) are a class of methods for building continuous-time normalizing flows in a simulation-free manner. These methods transform an arbitrary source distribution into a target distribution via a specified path of measures, which is modeled through a learned vector field. These methods have been extended to setting of Riemannian manifolds (Chen and Lipman, 2023) and infinite-dimensional spaces (Kerrigan et al., 2023).
Notably, these works do not approximate an optimal transport between the source and target measures and instead use an independent coupling, although stochastic interpolants (Albergo and Vanden-Eijnden, 2022; Albergo et al., 2023b) may be rectified (Liu et al., 2022; Lee et al., 2023) through an iterative optimization procedure. Importantly, the paths in flow matching (Lipman et al., 2022) maybe taken to be conditionally optimal, but marginally they are not. Pooladian et al. (2023a) and Tong et al. (2023) propose instead to couple the source and target distributions via optimal transport.
While some works (Lipman et al., 2022; Davtyan et al., 2023; Gebhard et al., 2023) have studied flow matching approaches for conditional distributions, these approaches simply provide additional conditioning information to the model and do not study COT. Concurrent work (Isobe et al., 2024) proposes extended flow matching, which learns a matrix field to perform conditional generation. However, the results of Isobe et al. (2024) are limited to the finite-dimensional setting and do not develop the theory of COT.
At the time of our submission, concurrent work by Barboni et al. (2024) and Chemseddine et al. (2024) appeared with similar results. Our results were developed independently from these works.
@@ 3 Background and Notation
Let $X,X^{\prime},X^{\prime\prime}$ represent arbitrary separable Hilbert spaces, equipped with the Borel $\sigma$-algebra. We use $\mathbb{P}(X)$ to represent the space of Borel probability measures on $X$, and $\mathbb{P}_{p}(X)\subseteq\mathbb{P}(X)$ to represent the subspace of measures having finite $p$th moment. If $\eta\in\mathbb{P}(X)$ is a probability measure on $X$ and $T:X\to X^{\prime}$ is measurable, then the pushforward measure $T_{\#}\eta({-})=\eta(T^{-1}({-}))$ is a probability measure on $X^{\prime}$. Maps of the form $\pi^{X}:X\times X^{\prime}\to X$ represent the canonical projection onto the respective space. We will also use maps such as $\pi^{1}:X\times X^{\prime}\times X^{\prime\prime}\to X$ and $\pi^{1,2}:X\times X^{\prime}\times X^{\prime\prime}\to X\times X^{\prime}$ to denote projection onto specific coordinates. The set $C_{b}(X)$ represents the space of continuous and bounded functions and $C^{\infty}_{c}(X)$ is the space of smooth and compactly supported functions.

### v2

# Dynamical Conditional Optimal Transport through Simulation-Free Flows

@@ Abstract
We study the geometry of conditional optimal transport (COT) and prove a dynamical formulation which generalizes the Benamou-Brenier Theorem. Equipped with these tools, we propose a simulation-free flow-based method for conditional generative modeling. Our method couples an arbitrary source distribution to a specified target distribution through a triangular COT plan, and a conditional generative model is obtained by approximating the geodesic path of measures induced by this COT plan. Our theory and methods are applicable in infinite-dimensional settings, making them well suited for a wide class of Bayesian inverse problems. Empirically, we demonstrate that our method is competitive on several challenging conditional generation tasks, including an infinite-dimensional inverse problem.
@@ 1 Introduction
Many fundamental tasks in machine learning and statistics may be posed as modeling a conditional distribution $\nu(u\mid y)$, but where obtaining an analytical representation of $\nu(u\mid y)$ is often impractical. While sampling-based approaches such as Markov Chain Monte Carlo (MCMC) methods are useful, they suffer from several limitations. First, MCMC requires numerous likelihood evaluations, rendering it prohibitively expensive in scientific and engineering applications where the likelihood is determined by an expensive numerical simulator. Second, MCMC must be run anew for every observation $y$, which is impractical in applications such as Bayesian inverse problems (Dashti and Stuart, 2013) and generative modeling (Mirza and Osindero, 2014). These limitations motivate the need for a likelihood-free (Cranmer et al., 2020) and amortized (Amos et al., 2023) approach. While methods like ABC (Beaumont, 2010) and variational inference (Blei et al., 2017) address these challenges, they are difficult to scale to high dimensions or have limited flexibility.
Recently, generative models such as normalizing flows (Papamakarios et al., 2019, 2021), GANs (Ramesh et al., 2022), and diffusion models (Sharrock et al., 2022) have shown promise in amortized and likelihood-free inference. These models may be viewed in the framework of measure transport (Baptista et al., 2020), where samples $u\sim\eta(u)$ from a tractable source distribution are transformed by a mapping $T(y,u)$ such that the transformed samples are approximately distributed as $\nu(u\mid y)$. One way to achieve this is through triangular mappings (Baptista et al., 2020; Spantini et al., 2022), where a joint source distribution $\eta(y,u)$ is transformed by a mapping of the form $T:(y,u)\mapsto(T_{Y}(y),T_{U}(y,u))$. Under suitable assumptions, if $T$ transforms the source $\eta(y,u)$ into the target $\nu(y,u)$, then $T_{U}(y,{-})$ couples the conditionals $\eta(u\mid y)$ and $\nu(u\mid y)$.
Typically, such a map $T$ is not unique (Wang et al., 2023), and a natural idea is thus to regularize the transport and search for an admissible mapping that is in some sense optimal. In other words, learning a conditional sampler may be phrased as finding a conditional optimal transport (COT) map. While there exists some work on learning COT maps, these approaches often rely on a difficult adversarial optimization problem (Baptista et al., 2020; Hosseini et al., 2023; Bunne et al., 2022; Ray et al., 2022) or simulating from the model during training (Baptista et al., 2023; Wang et al., 2023). In this work, we propose a conditional generative model for likelihood-free inference based on a dynamic formulation of conditional optimal transport. Specifically, our contributions are as follows:
@@
    We develop a general theoretical framework for dynamic conditional optimal transport in separable Hilbert spaces. Our framework is applicable in infinite-dimensional spaces, enabling applications in function space Bayesian inference. In Section 4, we study the conditional Wasserstein space $\mathbb{P}_{p}^{\mu}(Y\times U)$ and show that this space admits constant speed geodesics between any two measures. In Section 5, we characterize the absolutely continuous curves of measures in $\mathbb{P}_{p}^{\mu}(Y\times U)$ via the continuity equation and triangular vector fields. As a consequence, we obtain conditional generalizations of the McCann interpolants (McCann, 1997) and the Benamou-Brenier Theorem (Benamou and Brenier, 2000).
@@
    In Section 6, we propose COT flow matching (COT-FM), a simulation-free flow-based model for conditional generation. This model directly leverages our theoretical framework, where we learn to model a path of measures interpolating between an arbitrary source and target distribution via a geodesic in the conditional Wasserstein space.
@@
    In Section 7, we demonstrate our method on several challenging conditional generation tasks. We apply our method to two Bayesian inverse problems – one arising from the Lotka-Volterra dynamical system, and an infinite-dimensional problem arising from the Darcy Flow PDE. Our method shows competitive performance against recent COT methods.
@@ 2 Related Work
Conditional Optimal Transport.
Conditional Optimal Transport (COT) remains relatively under-explored in both machine learning and related fields. Recent approaches learn static COT maps via input convex networks (Bunne et al., 2022; Wang et al., 2023) or normalizing flows (Wang et al., 2023). In addition, there have been a number of heuristic approaches to conditional simulation through W-GANs (Sajjadi et al., 2017; Adler and Öktem, 2018; Kim et al., 2022, 2023), for which Chemseddine et al. (2023) provide a rigorous basis. Closely related to our method are those which employ triangular plans (Carlier et al., 2016; Trigila and Tabak, 2016), which have been modeled through GANs in Euclidean spaces (Baptista et al., 2020) and function spaces (Hosseini et al., 2023). In contrast, our work uses a novel dynamic formulation of COT, which we model through a generalization of flow matching (Lipman et al., 2022; Albergo et al., 2023b). This allows us to use flexible architectures while avoiding the difficulties of training GANs (Arora et al., 2018).
@@
Flow matching (Lipman et al., 2022) and stochastic interpolants (Albergo et al., 2023a) are a class of methods for building continuous-time normalizing flows in a simulation-free manner. Notably, these works do not approximate an optimal transport between the source and target measures. Pooladian et al. (2023) and Tong et al. (2023) propose instead to couple the source and target distributions via optimal transport, leading to marginally optimal paths. In this work, we study an extension of these techniques for conditional generation.
While some works (Davtyan et al., 2023; Gebhard et al., 2023; Isobe et al., 2024; Wildberger et al., 2024) have applied flow matching for conditional generation, these approaches do not employ COT. Notably, the aforementioned approaches are limited to the finite-dimensional setting, whereas our method adds to the growing literature on function-space generative models (Hosseini et al., 2023; Kerrigan et al., 2023, 2024; Lim et al., 2023; Franzese et al., 2024). At the time of our submission, concurrent work by Barboni et al. (2024) and Chemseddine et al. (2024) appeared with similar results. Our results were developed independently from these works.
@@ 3 Background and Notation
Let $X,X^{\prime}$ represent arbitrary separable Hilbert spaces, equipped with the Borel $\sigma$-algebra. We use $\mathbb{P}(X)$ to represent the space of Borel probability measures on $X$, and $\mathbb{P}_{p}(X)\subseteq\mathbb{P}(X)$ to represent the subspace of measures having finite $p$th moment. If $\eta\in\mathbb{P}(X)$ is a probability measure on $X$ and $T:X\to X^{\prime}$ is measurable, then the pushforward measure $T_{\#}\eta({-})=\eta(T^{-1}({-}))$ is a probability measure on $X^{\prime}$. Maps of the form e.g. $\pi^{X}:X\times X^{\prime}\to X$ represent the canonical projection.